OpenStudy (anonymous):

Two iron weights, one twice the mass of the other, are dropped from the top of a building. Compared with the lighter weight, the twice as heavy weight will hit the ground with

OpenStudy (anonymous):

I believe they will hit with the same momentum since the momentum of a falling (gravity) object towards earth is not dependent on mass.

OpenStudy (anonymous):

I'm not sure what he was quantity you are asking for. If you are asking for the speed of the two objects, then the speed will be the same for both objects. This is because gravity accelerates all objects at the same rate g, and so both objects will have speed v = g * t ; where t is how long the objects fell for. However, if momentum is what you are asking for, then the momentum is NOT the same. Momentum is defined as: p = m * v The speeds for the objects are the same, but one mass is twice the other! p1 = m * v p2 = 2m * v So the heavier object hits the ground with 2x the momentum as the lighter one.

OpenStudy (anonymous):

Correct answer: twice the momentum